OpenStudy (anonymous):

Can someone guide me through the steps of (3x-x^2)/(2x-2)? I understand the concept of polynomial long division, I am just confused on how to deal with the coefficient 2 on the 2x in the denominator.

OpenStudy (anonymous):



OpenStudy (anonymous):

\[2(2-x+\frac{ 2 }{ x-1 })\]

OpenStudy (anonymous):

Why is the 2 taken out of 2x-1?

OpenStudy (anonymous):

It becomes simpler to solve

OpenStudy (anonymous):

Oh... so you temporarily take it out, do the math, then add it back in to the final answer.

OpenStudy (anonymous):

exactly, a little tricky

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

Oh and sorry it should be 1/2 not 2 since it was on the denominator

OpenStudy (anonymous):

so final equation is actually \[\frac{ 1 }{ 2 }(2-x+\frac{ 2 }{ x-1 })\]